OpenStudy (anonymous):

Suppose cos θ = – 0.418 and 180º < θ < 270º. Find sin θ.

OpenStudy (apoorvk):

\[sin(\theta) = \sqrt{1-cos^2(\theta)}\] Now, since its provided that theta lies in the third quadrant, and sin(theta) is negative in the third quadrant, don't forget to prefix a minus sugn before the value of sin(theta) that you get.

OpenStudy (anonymous):

is cos^2(x) the same as cos(x)^2?

OpenStudy (apoorvk):